# Enhanced thermoelectric effects in a driven one-dimensional system

### v1

@@ Abstract
We investigate the thermoelectric properties of a one-dimensional quantum system in the presence of an external driving. We employ Floquet scattering theory to calculate linear-response stationary thermoelectric figures of merit in a single-channel conductor subjected to a periodically varying delta-like potential barrier. We include a step barrier in one of the leads as a model of a nanoscale inhomogeneous semiconducting system. In the absence of a step barrier, we found that the external driving can strongly enhance the Seebeck coefficient, particularly at low temperatures, with a relative increase with respect to the static condition reaching up to 200% at large frequencies. In the presence of a step barrier, we found that the thermoelectric Onsager coefficient for the driven case is also enhanced with respect to the static case with a significant effect at low temperatures when the chemical potential is within the gap of the semiconductor. Our results demonstrate that external driving can be used to tune and enhance the thermoelectric capabilities of low-electron-density nanodevices.
@@ I Introduction
Heat management is key to achieving high qubit density and good performance in quantum processors [1]. Thermoelectric effects provide us with a clever way to cooling via optimally controlling heat flows and, at the same time, convert it into electricity [2], which can be reused at the processor level or easily piped to higher temperatures where more efficient cooling mechanisms can be employed [3]. In this perspective, the thermoelectric properties of mesoscopic nanodevices [4] routinely employed in quantum technologies are of immediate interest. A paradigmatic case are thermoelectric effects in superconductors, which are at the core of several recently proposed thermoelectrically based quantum devices [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
Operating nano-devices through external periodic AC driving (or photon-assisted mechanisms) is a natural approach to heat management (for example for the implementation of heat engines, refrigerators and heat pumps) and for improving their thermoelectric capabilities, possibly taking advantage of quantum effects, see for example Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. The vast majority of such papers deals with the adiabatic regime, where the frequency of the periodic driving is the smallest energy scale, while only a few concern high frequencies, beyond the adiabatic regime. A photon-assisted refrigerator has been experimentally realized in Ref. [58].
@@
From the modeling point of view, quantum coherent transport for non-interacting periodically driven systems can be fruitfully described using the Floquet scattering theory [59, 60, 61], which can be considered as the analogue of the Bloch theory in the time domain.
@@
In this paper, we investigate the effect of AC driving on the stationary linear thermoelectric current in one-dimensional, quantum-coherent, non-interacting electronic systems. Specifically, we employ the Floquet scattering formalism to calculate the electric and thermoelectric Onsager coefficients for a system consisting of a driven tunnel barrier located between two leads. This analysis also accounts for an energy band shift between the leads, modeled as a step potential. We study the Onsager and Seebeck coefficients as functions of temperature and driving frequency, considering various parameters such as driving amplitude, the strength of the static barrier component, and the height of the step potential. In particular, we address the high-frequency regime, where the driving frequency can be comparable to other energy scales, including the chemical potential, temperature, or energy gaps. This regime is experimentally accessible in systems with sufficiently low electron density, such as semiconductors or two-dimensional materials, where electron density can be readily controlled by gating. We find that driving can enhance the Seebeck coefficient, particularly at low temperatures, by increasing the strength of the AC component of the barrier. Similarly, in the presence of an energy band shift, we observe an enhancement of the thermoelectric Onsager coefficient compared to the corresponding static case.
@@ II Floquet transport
[…]
where $E_{n}=E+n\hbar\omega$, with $n\in\mathbb{Z}$. The integration over the energies and the summation $\sum_{n}^{{}^{\prime}}$ are computed only for propagating states in the leads, i.e., $E>\epsilon_{\alpha}$ and $E_{n}>\epsilon_{\beta}$, where $\epsilon_{i}$ denotes the energy band bottom for $i$-lead. In Eq. (1), $S(E_{2},E_{1})_{\alpha\beta}$ is the scattering amplitude for an electron injected in lead $\beta$ with energy $E_{1}$ and arriving in lead $\alpha$ with energy $E_{2}$, and $f_{\alpha}(E)=\{1+\exp[(E-\mu_{\alpha})/k_{B}T_{\alpha}]\}^{-1}$ is the Fermi function relative to lead $\alpha$, where $\mu_{\alpha}$ and $T_{\alpha}$ are, respectively, its temperature and chemical potential. Thus $\left|S(E,E_{n})_{\alpha\beta}\right|^{2}$ represents the probability for the transfer of an electron from lead $\beta$ to $\alpha$ with the emission of $n$ photons of energy $\hbar\omega$. For the propagating states, one can show the unitarity conditions [59]
$\sum_{n,\beta}\,^{{}^{\prime}}|S(E,E_{n})_{\alpha\beta}|^{2}=\sum_{n,\alpha}\,^{{}^{\prime}}|S(E_{n},E)_{\alpha\beta}|^{2}=1\ .$ (2)
@@
[…]
@@
$\mathcal{T}_{\alpha\beta}(E)=\sum_{n}\,^{{}^{\prime}}\left|S(E_{n},E)_{\alpha\beta}\right|^{2},$ (4)
@@
Since we will investigate how the AC driving affects the thermoelectric quantities in linear response, it is useful to introduce the linear expansion of the current for a system at temperature $T$ in terms of the affinities. Specializing to a two terminal device we can write the affinities as $\mathcal{F}_{e}=\Delta V/T$ and $\mathcal{F}_{h}=\Delta T/T^{2}$ for small voltage bias $\Delta V\ll k_{B}T/e$ and temperature difference $\Delta T\ll T$. We also assume that chemical potentials and temperatures are applied symmetrically, i.e., $\mu_{L/R}=\mu\pm e\Delta V/2$ and $T_{L/R}=T\pm\Delta T/2$. Thus the current can be written as
@@
where $J_{e}^{P}$ is the pumped current which is zero when $\mathcal{T}_{RL}(E)=\mathcal{T}_{LR}(E)$, see App. B. The Onsager coefficients $L_{\rm ee}$ and $L_{\rm eh}$ can be directly connected to relevant physical quantities such as the two terminal conductance $G=T^{-1}L_{\rm ee}$ and the Seebeck coefficient [63]
@@
and
@@
where we define $\bar{\mathcal{T}}(E)=[\mathcal{T}_{RL}(E)+\mathcal{T}_{LR}(E)]/2$ and $f^{\prime}(E)=\partial f/\partial E$ is the energy derivative of the Fermi function $f(E)$, whose chemical potential and temperature are $\mu$ and $T$, respectively. This shows that the expressions of these two Onsager coefficients are formally equal to the ones relative to the stationary case, where $\bar{\mathcal{T}}(E)$ is replaced by the transmission in the stationary case  [4]. Due to the even symmetry in energy of the Fermi function derivative, $L_{\rm ee}$ is determined by the even-in-energy component of $\bar{\mathcal{T}}(E)$ around the Fermi energy, while $L_{\rm eh}$ measures the odd-in-energy symmetry of $\bar{\mathcal{T}}(E)$ around the Fermi energy.
Furthermore, we note that for very small temperatures we can expand in energy the transmission function around the chemical potential such as $\bar{\mathcal{T}}(E)=\bar{\mathcal{T}}(\mu)+\bar{\mathcal{T}}^{\prime}(\mu)(E-\mu)+\mathcal{O}[(E-\mu)^{2}]$. Using the Sommerfeld expansion one can show that $L_{\rm ee}\simeq e^{2}\bar{\mathcal{T}}(\mu)T/h$ and $L_{\rm eh}\simeq(\pi^{2}/3)(e/h)k_{B}^{2}\bar{\mathcal{T}}^{\prime}(\mu)T^{3}$ [4]. As a consequence, in the same limit one finds that $G$ is independent of temperature, while the Seebeck coefficient is linear in temperature, namely
where the dimensions are given by the prefactor $k_{B}/e\simeq 81\mu$V/K.
III Model
In this section we apply the Floquet approach to an ideal one-dimensional setup consisting of a single band electronic system with a driven time-periodic delta-like impurity. The system is described by the following Hamiltonian
$H=\frac{\hbar^{2}k^{2}}{2m}+W(t)\delta(x)+\Delta\theta(x),$ (10)
where $W(t)=W_{0}+2W_{1}\cos(\omega t)$ is the time-dependent strength of a delta-like potential. Notice that the potential is a periodic function, i.e. $W(t+T)=W(t)$ with period $T=2\pi/\omega$. Moreover, we also include the possibility of an energy band shift $\Delta=\epsilon_{R}-\epsilon_{L}$, which represents the height of a band-mismatch located in $x=0$ ($\epsilon_{R}$ and $\epsilon_{L}$ are the energies band bottom of the right lead and left lead, respectively). $\theta(x)$ is the Heaviside function, see Fig. 1]. Note that we consider a single driving parameter so that we cannot realize adiabatic pumping.
$x$$E$$\mu$$W(t)\delta(x)$$\Delta$
@@ III Model
where $k_{n}=\sqrt{2mE_{n}}/\hbar$ and $k^{\prime}_{n}=\sqrt{2m(E_{n}-\Delta)}/\hbar$ are the wavevectors in the left and right leads, respectively, relative to the energy $E_{n}=E+n\hbar\omega$. The term with $a_{n}$ represents right-going incident electrons in the left lead. By choosing $a_{n}=\delta_{n,0}$ we set their energy to $E_{0}=E$ and we normalize the flux to 1. The coefficient $\rho_{n}$ represents the left-to-left reflection amplitude, i.e. $\rho_{n}=S_{LL}(E_{n},E)$, for an incoming electron with energy $E$ to be scattered back at energy $E_{n}$. When $n\neq 0$ it corresponds to a process of inelastic reflection where the final propagating energy state $E_{n}>0$ has energies higher ($n>0$) or lower ($n<0$) than the initial state due to the interaction with the driven delta-like barrier. Analogously, the coefficient $\xi_{n}$ represents the left-to-right transmission amplitude $\xi_{n}=S_{RL}(E_{n},E)$, non-zero only if $E_{n}>\Delta$. From the scattering matrices we can similarly define the total transmission functions according to Eq. (4). Notice that in the case of time-independent potential, we have a finite contribution only from the term with $n=0$.
@@
Equation (13) consists of an infinite set of linear equations for the coefficients $\xi_{n}$. It can be conveniently recast in a matrix form as $\mathbf{M\cdot}\vec{\xi}=\vec{B}$, with $\mathbf{M}$ a matrix with infinite dimensions, and infinite dimensional vectors $\vec{\xi}=(...,\xi_{1},\xi_{0},\xi_{-1},...)$ and $\vec{B}=(...,b_{1},b_{0},b_{-1},...)$, where $b_{n}=-ik_{n}a_{n}$. For a given $\vec{B}$, which is fixed by the given incident coefficients $a_{n}$, such a matrix equation must be solved for the vector $\vec{\xi}$. Due to the infinite dimensionality of the matrix $\mathbf{M}$, it is hard in practice to solve it by direct inversion $\vec{\xi}=\mathbf{M}^{-1}\cdot\vec{B}$. However, in the case of weak driving potentials, it is convenient to use a perturbative approach in the small parameter $v=W_{1}/W_{0}\ll 1$, where one can effectively truncate the matrix $\mathbf{M}$ to a finite dimension. For example, for sufficiently small $v$, one can solve the problem considering only the energy $E$ and the first two sidebands $E_{\pm 1}=E\pm\hbar\omega$, which corresponds to the $3\times 3$ matrix
@@
While in this case $\mathbf{M}^{-1}$ can be computed analytically (see App. A), for a larger dimension it can be calculated numerically. In particular, to reach $n$-th order accuracy, we need a $12n+$ 1 dimensional matrix $\mathbf{M}$. To achieve higher accuracy, we just need to increase the dimension of $\mathbf{M}$ until convergence is attained. In the present work, for our numerical results we compute them with an accuracy of at least $10^{-7}$.
@@ IV.1 Single barrier
Figure 2: Total transmission $\mathcal{T}(E)$ as the function of energy $E$. The black curves are for the case without AC signal, i.e. $v=0$ and serves as a reference. (a) Comparison between the transmission $\mathcal{T}(E)$ with $\eta=0.1$ (solid) and $\eta=0.4$ (dashed). The colored curves correspond to different driving frequencies $\hbar\omega/\mu=0.2$ (blue), $0.5$ (red) and $0.9$ (green) and the vertical dotted lines are guides for the eyes, indicating the corresponding frequencies of the AC signal. The thin gray vertical line indicates the chemical potential ($E=\mu$). The inset is the zoom-in for $\eta=0.4$ around the chemical potential where $\hbar\omega/\mu=0.5$ (red) and $0.2$ (blue). (b) Zoom-in for the kink structure around $E-\mu=-0.8\mu$ for $\eta=0.1$. Curves with different colors correspond to $v=0.2$ (red), $0.3$ (orange), $0.4$ (blue).
@@
In Fig. 2(a) we assess the role of $\omega$ on the total transmission, for a fixed value of $v=0.4$ and two different values of $\eta=0.1$ (solid lines),0.4,0.4, 0.4 (dashed lines). The colored (solid and dashed) curves correspond to different values of the driving frequencies $\omega$ (see caption). The plots show that the main feature produced by the AC driving is a kink structure corresponding to $E=\hbar\omega$ and its multiples, the details of which depend on the parameters. For $\eta=0.1$, for example, the plots in panel (a) show that $\mathcal{T}(E)$ closely follows the black dashed curve up to $E=\hbar\omega$, and after the kink, it drops slightly. The effect appears to be stronger for $\hbar\omega=0.2\mu$ (blue curve). Note that in the whole range of energies considered, the value of the transmission is close to 1. When $\eta=0.4$, for a fixed strength of the static delta-like potential, the total transmission reduces in the whole range of energies, as shown in the same panel with dashed lines. However, at the same time, the relative strength of the kinks with respect the overall transmission profile increases. In general, the curves corresponding to the AC driving deviate more markedly from the static case (black curve), and the kink structures become more pronounced for increasing driving frequencies, such as the red and green curves. Notice that for energies $E\gg\hbar\omega$ all the curves with AC driving tend to the same curve, eventually converging from below to the curve for the static case. The origin and the shape of such kink structures is discussed in detail, using an analytical perturbative approach, in App. A.
As already mentioned, the AC driving induces subtle higher harmonic effects that extend beyond the fundamental driving frequency. Specifically, the transmission function $\mathcal{T}(E)$ exhibits modifications not only around the primary driving energy $E=\hbar\omega$, but also at integer multiples $E=n\hbar\omega$, where $n$ represents the harmonic order. However, the effect at higher harmonics is strongly suppressed, as it is higher order in the perturbation parameter $v$. In the inset of Fig. 2(a), a detailed zoom around $E=\mu$, corresponding to $E=2\hbar\omega$, reveals a particularly intriguing feature for the red dashed curve. Namely, even a weak kink structure emerging very close to the Fermi energy could lead to measurable consequences in the Seebeck coefficient (see below).
The role of the AC driving amplitude $v$ on the total transmission is investigated in panel (b) of Fig. 2. Results are presented for $v=0.4$ (blue curve),0.3,0.3, 0.3 (orange curve),0.2,0.2, 0.2 (red curve), 0 (black curve), at a fixed frequency $\hbar\omega=0.2\mu$ and $\eta=0.1$. With increasing $v$, the influence of higher harmonics of the driving frequency also emerge (not shown).

### v2

@@ Abstract
We investigate the thermoelectric properties of a one-dimensional quantum system in the presence of an external driving. We employ Floquet scattering theory to calculate linear-response stationary thermoelectric figures of merit in a single-channel conductor subjected to a periodically varying delta-like potential barrier. We also include a step barrier in one of the leads as a model of a nanoscale inhomogeneous semiconducting system. In the absence of a step barrier, we found that external driving can significantly enhance the Seebeck coefficient, particularly at low temperatures, with a relative increase of up to 200% at high frequencies compared to the static case. In the presence of a step barrier, we found that the thermoelectric Onsager coefficient for the driven case is also enhanced compared to the static case, with a significant photon-assisted effect at low temperatures when the chemical potential is within the semiconductor’s gap. Our results demonstrate that external driving can be used to tune and enhance the thermoelectric capabilities of low-electron-density nanodevices.
@@ I Introduction
Heat management is crucial for achieving high qubit density and optimal performance in quantum processors [1]. Thermoelectric effects provide us with a clever way to cool via optimally controlling heat flows and, at the same time, convert it into electricity [2], which can be reused at the processor level or easily piped to higher temperatures where more efficient cooling mechanisms can be employed [3]. In this perspective, the thermoelectric properties of mesoscopic nanodevices [4] routinely employed in quantum technologies are of immediate interest. A paradigmatic case are thermoelectric effects in superconductors, which are at the core of several recently proposed thermoelectrically based quantum devices [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
Operating nano-devices through external periodic AC driving (or photon-assisted mechanisms) is a natural approach to heat management (for example, for the implementation of heat engines, refrigerators, and heat pumps) and for improving their thermoelectric capabilities, possibly taking advantage of quantum effects, see for example Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. The vast majority of such papers deal with the adiabatic regime, where the frequency of the periodic driving is the smallest energy scale, whereas only a few concern high frequencies, which lie beyond the adiabatic regime. A photon-assisted refrigerator has been experimentally realized in Ref. [59].
@@
From a modeling perspective, quantum coherent transport for non-interacting periodically driven systems can be fruitfully described using Floquet scattering theory [60, 61, 62, 63], which can be considered an analogue of the Bloch theory in the time domain.
@@
In this paper, we investigate the effect of AC driving on the stationary linear thermoelectric current in one-dimensional, quantum-coherent, non-interacting electronic systems. Specifically, we employ the Floquet scattering formalism to calculate the electric and thermoelectric Onsager coefficients for a system comprising a driven tunnel barrier situated between two leads. This analysis also accounts for an energy band shift between the leads, modeled as a step potential. We study the Onsager and Seebeck coefficients as functions of temperature and driving frequency, considering various parameters, including driving amplitude, the strength of the static barrier component, and the height of the step potential. In particular, we focus on the high-frequency regime, where the driving frequency can be comparable to other energy scales, such as the chemical potential, temperature, or energy gaps. This regime is experimentally accessible in systems with sufficiently low electron density, such as semiconductors or two-dimensional materials, where electron density can be readily controlled by gating. We find that driving can enhance the Seebeck coefficient, particularly at low temperatures, by increasing the strength of the AC component of the barrier. Similarly, in the presence of an energy band shift, we observe an enhancement of the thermoelectric Onsager coefficient compared to the corresponding static case.
@@ II Floquet transport
$J_{e}^{\alpha}=\frac{e}{h}\!\!\int_{-\infty}^{\infty}\!\!\!\!\!\!\!\!dE\sum_{\beta,n}\,^{{}^{\prime}}\left|{\cal S}_{\alpha\beta}(E,E_{n})\right|^{2}\left[f_{\beta}(E_{n})-f_{\alpha}(E)\right],$ (1)
where $E_{n}=E+n\hbar\omega$, with $n\in\mathbb{Z}$. The integration over the energies and the summation $\sum_{n}^{{}^{\prime}}$ are computed only for propagating states in the leads, i.e., $E>\epsilon_{\alpha}$ and $E_{n}>\epsilon_{\beta}$, where $\epsilon_{i}$ denotes the energy band bottom for $i$-lead. In Eq. (1), ${\cal S}_{\alpha\beta}(E_{2},E_{1})$ is the scattering amplitude for an electron injected from lead $\beta$ toward the scattering region with energy $E_{1}$ and arriving in lead $\alpha$ with energy $E_{2}$, and $f_{\alpha}(E)=\{1+\exp[(E-\mu_{\alpha})/k_{B}T_{\alpha}]\}^{-1}$ is the Fermi function relative to lead $\alpha$, where $\mu_{\alpha}$ and $T_{\alpha}$ are, respectively, its temperature and chemical potential. Thus $\left|{\cal S}_{\alpha\beta}(E,E_{n})\right|^{2}$ represents the probability for the transfer of an electron from lead $\beta$ to $\alpha$ with the emission of $n$ photons of energy $\hbar\omega$. For the propagating states, one can show the unitarity conditions [60]
$\sum_{n,\beta}\,^{{}^{\prime}}|{\cal S}_{\alpha\beta}(E,E_{n})|^{2}=\sum_{n,\alpha}\,^{{}^{\prime}}|{\cal S}_{\alpha\beta}(E_{n},E)|^{2}=1\ .$ (2)
@@
$J_{e}^{\alpha}=\frac{e}{h}\int_{-\infty}^{\infty}\!\!\!\!\!dE\sum_{\beta}\left[\mathcal{T}_{\alpha\beta}(E)f_{\beta}(E)-\mathcal{T}_{\beta\alpha}(E)f_{\alpha}(E)\right].$ (3)
@@
$\mathcal{T}_{\alpha\beta}(E)=\sum_{n}\,^{{}^{\prime}}\left|{\cal S}_{\alpha\beta}(E_{n},E)\right|^{2},$ (4)
@@
Since we will investigate how the AC driving affects the thermoelectric quantities in linear response, it is useful to introduce the linear expansion of the current for a system at temperature $T$ in terms of the affinities. Specializing to a two-terminal device, we can write the affinities as $\mathcal{F}_{e}=\Delta V/T$ and $\mathcal{F}_{h}=\Delta T/T^{2}$ for small voltage bias $\Delta V\ll k_{B}T/e$ and temperature difference $\Delta T\ll T$. We also assume that chemical potentials and temperatures are applied symmetrically, i.e., $\mu_{L/R}=\mu\pm e\Delta V/2$ and $T_{L/R}=T\pm\Delta T/2$. Thus, the current can be written as
@@
where $J_{e}^{P}$ is the pumped current which is zero when $\mathcal{T}_{RL}(E)=\mathcal{T}_{LR}(E)$, see App. B. The Onsager coefficients $L_{\rm ee}$ and $L_{\rm eh}$ can be directly connected to relevant physical quantities such as the two-terminal conductance $G=T^{-1}L_{\rm ee}$ and the Seebeck coefficient [65]
@@
$L_{\rm ee}$ $=-\frac{e^{2}}{h}T\int dE\bar{\mathcal{T}}(E)f^{\prime}(E),$ (7)
and
$L_{\rm eh}$ $=-\frac{e}{h}T\int dE(E-\mu)\bar{\mathcal{T}}(E)f^{\prime}(E),$ (8)
where we define $\bar{\mathcal{T}}(E)=[\mathcal{T}_{RL}(E)+\mathcal{T}_{LR}(E)]/2$ and $f^{\prime}(E)=\partial f/\partial E$ is the energy derivative of the Fermi function $f(E)$, whose chemical potential and temperature are $\mu$ and $T$, respectively. This shows that the expressions of these two Onsager coefficients are formally equal to the ones relative to the static case, where $\bar{\mathcal{T}}(E)$ is replaced by the transmission in the static case  [4]. Due to the even symmetry in energy of the Fermi function derivative, $L_{\rm ee}$ is determined by the even-in-energy component of $\bar{\mathcal{T}}(E)$ around the Fermi energy, while $L_{\rm eh}$ measures the odd-in-energy symmetry of $\bar{\mathcal{T}}(E)$ around the Fermi energy.
Furthermore, we note that for very small temperatures we can expand in energy the transmission function around the chemical potential such as $\bar{\mathcal{T}}(E)=\bar{\mathcal{T}}(\mu)+\bar{\mathcal{T}}^{\prime}(\mu)(E-\mu)+\mathcal{O}[(E-\mu)^{2}]$. Using the Sommerfeld expansion one can show that $L_{\rm ee}\simeq e^{2}\bar{\mathcal{T}}(\mu)T/h$ and $L_{\rm eh}\simeq(\pi^{2}/3)(e/h)k_{B}^{2}\bar{\mathcal{T}}^{\prime}(\mu)T^{3}$ [4]. As a consequence, in the same limit, one finds that $G$ is independent of temperature, while the Seebeck coefficient is linear in temperature, namely
@@
III Model
In this section, we apply the Floquet approach to an ideal one-dimensional setup consisting of a single band electronic system with a driven time-periodic delta-like impurity. The system is described by the following Hamiltonian
$H=\frac{\hbar^{2}k^{2}}{2m}+W(t)\delta(x)+\Delta\theta(x),$ (10)
where $W(t)=W_{0}+2W_{1}\cos(\omega t)$ is the time-dependent strength of a delta-like potential, and $\theta(x)$ is the Heaviside function. Notice that the potential is a periodic function, i.e. $W(t+T)=W(t)$ with period $T=2\pi/\omega$. Moreover, we also include the possibility of an energy band shift $\Delta=\epsilon_{R}-\epsilon_{L}$, which represents the height of a band-mismatch located in $x=0$ ($\epsilon_{R}$ and $\epsilon_{L}$ are the energy band bottoms of the right lead and left lead, respectively, see Fig. 1). Note that we consider a single driving parameter so that we cannot realize adiabatic pumping.
@@ III Model
where $k_{n}=\sqrt{2mE_{n}}/\hbar$ and $k^{\prime}_{n}=\sqrt{2m(E_{n}-\Delta)}/\hbar$ are the wavevectors in the left and right leads, respectively, relative to the energy $E_{n}=E+n\hbar\omega$. The term with $a_{n}$ represents right-going incident electrons in the left lead. By choosing $a_{n}=\delta_{n,0}$ we set their energy to $E_{0}=E$ and we normalize the flux to 1. The coefficient $\rho_{n}$ represents the left-to-left reflection amplitude, i.e. $\rho_{n}={\cal S}_{LL}(E_{n},E)$, for an incoming electron with energy $E$ to be scattered back at energy $E_{n}$. When $n\neq 0$, it corresponds to a process of inelastic reflection where the final propagating energy state $E_{n}>0$ has energies higher ($n>0$) or lower ($n<0$) than the initial state due to the interaction with the driven delta-like barrier. Analogously, the coefficient $\xi_{n}$ represents the left-to-right transmission amplitude $\xi_{n}={\cal S}_{RL}(E_{n},E)$, non-zero only if $E_{n}>\Delta$. The total transmission functions are defined according to Eq. (4). Notice that in the case of a time-independent potential, we have a finite contribution only from the term with $n=0$.
@@
[…]
@@
While in this case $\mathbf{M}^{-1}$ can be computed analytically (see App. A), for larger dimensions it can be calculated numerically. In particular, to reach $n$-th order accuracy, we need a $2n+1$ dimensional matrix $\mathbf{M}$. To achieve higher accuracy, we just need to increase the dimension of $\mathbf{M}$ until convergence is attained. In the present work, we compute our numerical results with an accuracy of at least $10^{-7}$.
@@ IV.1 Single barrier
Figure 2: Total transmission $\mathcal{T}(E)$ as the function of energy $E$. The black curves are for the case without AC signal, i.e. $v=0$ and serve as a reference. (a) Comparison between the transmission $\mathcal{T}(E)$ with $\eta=0.1$ (solid) and $\eta=0.4$ (dashed). The colored curves correspond to different driving frequencies $\hbar\omega/\mu=0.2$ (blue), $0.5$ (red) and $0.9$ (green) and the vertical dotted lines are guides for the eyes, indicating the corresponding frequencies of the AC signal. The thin gray vertical line indicates the chemical potential ($E=\mu$). The inset is the zoom-in for $\eta=0.4$ around the chemical potential where $\hbar\omega/\mu=0.5$ (red) and $0.2$ (blue). (b) Zoom-in on the kink structure around $E-\mu=-0.8\mu$ for $\eta=0.1$. Curves with different colors correspond to $v=0$ (black), $0.2$ (red), $0.3$ (orange), $0.4$ (blue).
@@
In Fig. 2(a) we assess the role of $\omega$ on the total transmission, for a fixed value of $v=0.4$ and two different values of $\eta=0.1$ (solid lines),0.4,0.4 (dashed lines). The colored (solid and dashed) curves correspond to different values of the driving frequencies $\omega$ (see caption). The plots show that the main feature produced by the AC driving is a kink structure corresponding to $E=\hbar\omega$ and its multiples, the details of which depend on the parameters. For $\eta=0.1$, for example, the plots in panel (a) show that $\mathcal{T}(E)$ closely follows the black dashed curve up to $E=\hbar\omega$, and after the kink, it drops slightly. The effect appears to be stronger for $\hbar\omega=0.2\mu$ (blue curve). Note that in the whole range of energies considered, the value of the transmission is close to 1. When $\eta=0.4$, for a fixed strength of the static delta-like potential, the total transmission reduces in the whole range of energies, as shown in the same panel with dashed lines. However, at the same time, the relative strength of the kinks with respect to the overall transmission profile increases. In general, the curves corresponding to the AC driving deviate more markedly from the static case (black curve), and the kink structures become more pronounced as the driving frequency increases, as seen in the red and green curves. Notice that for energies $E\gg\hbar\omega$ all the curves with AC driving tend to the same curve, eventually converging from below to the curve for the static case. The origin and the shape of such kink structures is discussed in detail, using an analytical perturbative approach, in App. A.
As already mentioned, the AC driving induces subtle higher harmonic effects that extend beyond the fundamental driving frequency. Specifically, the transmission function $\mathcal{T}(E)$ exhibits modifications not only around the primary driving energy $E=\hbar\omega$, but also at integer multiples $E=n\hbar\omega$, where $n$ represents the harmonic order. However, the effect at higher harmonics is strongly suppressed, as it is a higher order in the perturbation parameter $v$. In the inset of Fig. 2(a), a detailed zoom around $E=\mu$, corresponding to $E=2\hbar\omega$, reveals a particularly intriguing feature for the red dashed curve. Namely, even a weak kink structure emerging very close to the Fermi energy could lead to measurable consequences in the Seebeck coefficient (see below).
The role of the AC driving amplitude $v$ on the total transmission is investigated in panel (b) of Fig. 2. Results are presented for $v=0$ (black curve), $0.2$ (red curve), $0.3$ (orange curve), and $0.4$ (blue curve), at a fixed frequency $\hbar\omega=0.2\mu$ and $\eta=0.1$. With increasing $v$, the influence of higher harmonics of the driving frequency also emerge (not shown).